# The dual nature of GHZ9: coexisting AGN and star formation activity in a remote X-ray source at z=10.145

### v3

@@ Abstract
We present JWST/NIRSpec PRISM spectroscopic characterization of GHZ9 at z= 10.145 0.010, currently the most distant source detected by the Chandra X-ray Observatory. The spectrum reveals several UV high-ionization lines, including C II, Si IV, N IV], C IV, He II, O III], N III], and C III]. The prominent rest-frame equivalent widths (EW(C IV)65Å, EW(O III])28Å, EW(C III])48Å) show the presence of a hard AGN radiation field, while line ratio diagnostics are consistent with either AGN or star formation as the dominant ionizing source. GHZ9 is nitrogen-enriched (6–9.5 (N/O)⊙), carbon-poor (0.2–0.65 (C/O)⊙), metal-poor (Z = 0.01–0.1 Z⊙), and compact ( 106 pc), similarly to GNz11, GHZ2, and recently discovered N-enhanced high redshift objects. We exploited the newly available JWST/NIRSpec and NIRCam dataset to perform an independent analysis of the Chandra data confirming that GHZ9 is the most likely JWST source associated with X-ray emission at 0.5-7 keV. Assuming a spectral index = 2.3 (1.8), we estimate a black hole (BH) mass of 1.60 0.31 (0.48 0.09) 108M⊙, which is consistent either with Eddington-accretion onto heavy ( 106 M⊙) BH seeds formed at z=18, or super-Eddington accretion onto a light seed of 10 M⊙ at z = 25. The corresponding BH-to-stellar mass ratio MBH/Mstar= 0.330.22 (0.100.07), with a stringent limit 0.02, implies an accelerated growth of the BH mass with respect to the stellar mass. GHZ9 is the ideal target to constrain the early phases of AGN-galaxy coevolution with future multi-frequency observations.
@@ 1 Introduction
The James Webb Space Telescope (JWST) is revolutionizing our understanding of both galaxies and active galactic nuclei (AGN) in the high redshift Universe. Several surveys have found a density of bright galaxies at which is significantly larger than previously predicted (e.g., Castellano et al., 2022, 2023; Finkelstein et al., 2024; Chemerynska et al., 2024; McLeod et al., 2024). In addition, JWST NIRSpec (Jakobsen et al., 2022) observations have detected a higher-than-expected number of both Broad-Line AGN (BLAGN, e.g., Harikane et al., 2023; Maiolino et al., 2024a; Matthee et al., 2024) and Narrow-Line AGN (NLAGN, e.g., Chisholm et al., 2024; Curti et al., 2025) at high redshift, exceeding the predictions from extrapolated quasar (e.g., Niida et al., 2020; Shen et al., 2020) and AGN (e.g., Finkelstein & Bagley, 2022) luminosity functions, and from deep X-ray surveys (e.g., Giallongo et al., 2015, 2019).
@@
A comprehensive census of the AGN population at high redshifts is crucial for understanding the origin of the correlation between the physical properties of supermassive black holes (SMBHs) and their host galaxies, as found in the local Universe (e.g., Kormendy & Ho, 2013; Greene et al., 2020). In addition, a deeper understanding of the demographics and properties of distant AGN is necessary to assess their role in the reionization process (e.g., Giallongo et al., 2015; Madau et al., 2024; Dayal et al., 2025) and evaluate whether they contribute significantly to the ionizing UV emission from the bright galaxy population at z9, as suggested by recent studies (e.g., Maiolino et al., 2024a; Castellano et al., 2024; Harikane et al., 2025; Hegde et al., 2024). However, identifying the AGN population at high redshifts remains challenging.
@@
The demarcation lines between star-forming galaxies (SFGs) and AGN in the classic diagnostic diagrams (Baldwin et al., 1981; Veilleux & Osterbrock, 1987; Kauffmann et al., 2003) are less effective at high redshifts, with the two populations overlapping in the same regions (Kocevski et al., 2023; Maiolino et al., 2024a; Scholtz et al., 2023; Übler et al., 2023). New diagnostic diagrams have been calibrated (e.g., Calabrò et al., 2023; Hirschmann et al., 2023; Mazzolari et al., 2024b) and tested on a few z10 AGN candidates identified by JWST, such as GN-z11 (Maiolino et al., 2024b) and GHZ2 (Castellano et al., 2024).
@@
The discovery of AGN candidates beyond z=10, when the Universe was less than 450 Myr old, is pushing our exploration closer to the black hole seeding epoch, offering a unique opportunity to constrain the origin of SMBHs (Natarajan et al., 2024; Trinca et al., 2024b; Gordon et al., 2025; Hu et al., 2025; Jeon et al., 2025). As recently highlighted by Taylor et al. (2024), this remote epoch represents a unique window to probe black hole seeds, as by z 6, the black hole (BH) mass function has largely lost memory of its initial seeding phase (see also, Valiante et al., 2018).
@@
The SMBH masses associated with the Chandra X-ray detections of the z10 objects UHZ1 (Goulding et al., 2023; Bogdán et al., 2024), and GHZ9 (Kovács et al., 2024) have prompted discussions on alternative black hole seeding scenarios pathways beyond the standard light and heavy seed scenarios (Inayoshi et al., 2020), including models invoking accretion onto cosmological primordial black hole seeds (Dayal, 2024; Matteri et al., 2025; Zhang et al., 2025; Ziparo et al., 2025).
@@
In this paper, we present a detailed analysis of the physical properties of GHZ9, based on the JWST/NIRSpec PRISM data presented in Napolitano et al. (2025), which confirmed the object at z=10.145. We examine the contribution of AGN and star-formation to the spectrum of this source through rest-frame UV and optical diagnostic diagrams. We also take advantage of the newly available NIRSpec and NIRCam information on the sources in its vicinity to perform an independent analysis of the association with the Chandra X-ray emission, first presented by Kovács et al. (2024). GHZ9 provides a unique opportunity to test our understanding of AGN at high redshifts, as it is the most distant X-ray detected AGN known to date and benefits from several rest-frame optical and UV line detections thanks to JWST/NIRSpec.
@@ 2 NIRSpec observations and data analysis
GHZ9 (R.A. = 3.478756, Decl. = -30.345520) was identified as a high redshift candidate in the GLASS-JWST NIRCam field (Treu et al., 2022) by Castellano et al. (2023). It was observed using NIRSpec in the PRISM-CLEAR configuration as part of the Cycle 2 program GO-3073 (PI: Marco Castellano). The observation utilized three-shutter slits with a three-point nodding pattern for optimal background subtraction, with a total exposure time of 19701s over three separate visits.
@@
The detailed data reduction, as well as the GHZ9 spectrum and analysis (including redshift determination and line fitting), are presented in Napolitano et al. (2025). Briefly, data were processed using the standard calibration pipeline provided by STScI (version 1.13.4) and the Calibration Reference Data System (CRDS) mapping 1197, following the methodology of Arrabal Haro et al. (2023), which produces both 2D and 1D flux-calibrated spectra. To correct for potential slit-losses, the NIRSpec spectrum was calibrated against the most recent NIRCam broadband photometry (Merlin et al., 2024) by matching the continuum level. Additionally, since the source is magnified by the foreground Abell-2744 cluster, rest-frame quantities for GHZ9 were corrected for magnification ( = 1.36, Bergamini et al., 2023).
@@
The stellar masses of GHZ9 and of the two foreground interlopers ID = 29686 and ID = 29852 (see Fig. 1) were estimated using zphot (Fontana et al., 2000) as described by Santini et al. (2023), by fitting the observed HST and JWST photometry111The photometry of GHZ9 (ID = 29722) and of the two foreground sources is publicly available in the ASTRODEEP catalog (Merlin et al., 2024) https://vizier.cds.unistra.fr/viz-bin/VizieR?-source=J/A+A/691/A240 with Bruzual & Charlot (2003) templates, assuming delayed star-formation histories. The contribution from nebular continuum and line emission was included following Schaerer & de Barros (2009) and Castellano et al. (2014).
@@
We measured the half-light radius (re = 0.028′′) in the rest-frame UV using the same procedure adopted by Mascia et al. (2024), with the python software Galight222https://github.com/dartoon/galight (Ding et al., 2020). Assuming a Sérsic profile, with an axial ratio between –, and a Sérsic index of = 1, the fit was performed on the F150W NIRCam image. We visually inspected the result, finding no significant residuals in the luminosity profile.
@@
The spectroscopic redshift (zspec = 10.145 0.010) was determined from a weighted average of emission line centroids with S/N 5, calculated via direct integration. We measured the UV slope ( = -1.10 0.12) by fitting a power-law model () to the continuum flux at 1400–2600 Å rest-frame, after masking any potential emission features within the considered wavelength range. We employed emcee (Foreman-Mackey et al., 2013) for Markov chain Monte Carlo (MCMC) analysis.
@@
For emission lines with S/N 3, a Gaussian fit was applied to the continuum-subtracted flux using the specutils package from astropy (Astropy Collaboration et al., 2013) combined with emcee. Unresolved doublets and multiplets were modeled as single Gaussian profiles, while partially blended lines were fitted with double-Gaussian profiles (see Fig. A.1 from Napolitano et al., 2025, for visualization).
@@ 3 Chandra X-ray observations
We analyzed all 101 publicly available Chandra X-ray observations of Abell-2744 (Chadayammuri et al., 2024), excluding ObsId=2212, as it was taken with a different CCD (ACIS-S instead of ACIS-I). We reprocessed all of the observations in the whole 0.5–7 keV band where Chandra is sensitive using standard CIAO tools and created a mosaic, with the deepest part having an exposure of 2.14Ms. To ensure accurate astrometry, we matched sources detected in the 0.5–7 keV band to the Gaia DR2 catalog and found that a small translation of 0.27′′ was required, consistent with Kovács et al. (2024). GHZ9 has an average off-axis angle of 6.6′ from the deepest part of the X-ray field; the aperture radius including 90% of the encircled energy fraction (EEF) at 1.5 keV is 6′′. The net exposure time, accounting for vignetting, is 1.76 Ms. Given the presence of a nearby (4′′) bright X-ray star, we carefully accounted for its contamination.
@@
We extracted GHZ9 source counts from a circular region of 1′′ radius centered on the JWST position in the 0.5–7 keV band image, where a peak of emission is visible in X-rays (see Fig. 1). This region includes only 13% of the total source counts due to the EEF. To compute the total number of counts, we considered both the nearby star contamination at the location of GHZ9 and the background contribution, evaluated from a nearby source-free circular region of 100 arcsec2. We also verified that the background is stable over a much larger region. In particular, we found that the background at the position of GHZ9 is consistent with the average background measured at the position of other X-ray sources detected outside the region affected by the diffuse emission from the cluster, whose peak (outskirts) is at 5.5 (2.8) arcmin from GHZ9. We measured 35.96.9 counts in the 0.5–7 keV band for GHZ9.
@@
We have performed additional checks on the X-ray detection. As a first test, we ran the Chandra detection tool for point-like sources wavdetect in a 100100 arcsec2 region centered on GHZ9, after removing the star contamination. GHZ9 is detected with 23.27.6 counts in the observed-frame 0.5–7 keV band (3.3 significance). The flux and luminosity values discussed in the following would be scaled down by 35%. As a second check, we evaluated whether the shape of the X-ray continuum of GHZ9 differs from that of the nearby star by computing the ratio between the counts in the soft (S=0.5–2 keV) and hard (H=2–7 keV) bands for both objects. We found this test to be inconclusive due to the large uncertainties resulting in S/H=2.0 for GHZ9 and 1.2 for the star. As a final check, we divided the Chandra dataset into two subsets with comparable exposure time. Unsurprisingly, the source is formally undetected in each individual subset. However, the counts measured within a 1″ radius circular region in both subsets are consistent with each other and with the expected signal given the reduced exposure time, thus excluding the possibility that the detection is due to a spurious fluctuation in the counts.
@@
The 0.5–7 keV counts for GHZ9 were then converted into flux and luminosity (both de-magnified) assuming a power-law model with two possible spectral indexes =1.8 (Piconcelli et al., 2005) and =2.3 (as in Kovács et al., 2024), which is comparable to the values measured in luminous quasars at z6–7.5 (Zappacosta et al., 2023). The assumption of the power-law, related to the main AGN emission process in X-rays (inverse Compton emission of accretion disk photons with energetic electrons in the corona), is required to transform the source count rate, which is a detector-dependent measure, into physical quantities, like flux and luminosity. In the conversion process, we have considered the Chandra effective area corresponding to Cycle 24, during which most of the Abell 2744 observations were carried out. The observed-frame 2–10 keV flux is 1.310-16 erg cm-2s-1 if =2.3 (2.010-16 erg cm-2s-1 if =1.8). The corresponding rest-frame 2–10 keV luminosity is 3.81044 erg s-1 (1.81044 erg s-1). For the observed-frame 0.5–3 keV band, the flux measured, 2.210-16 erg cm-2s-1, is consistent with the value reported by Kovács et al. (2024), assuming the same photon index and magnification factor.

### v4

# The Dual Nature of GHZ9: Coexisting Active Galactic Nuclei and Star Formation Activity in a Remote X-Ray Source at z=10.145

@@ Abstract
We present James Webb Space Telescope (JWST)/NIRSpec PRISM spectroscopic characterization of GHZ9 at z= 10.145 0.010, currently the most distant source detected by the Chandra X-ray Observatory. The spectrum reveals several UV high-ionization lines, including C II, Si IV, N IV], C IV, He II, O III], N III], and C III]. The prominent rest-frame equivalent widths (EW(C IV)65Å, EW(O III])28Å, EW(C III])48Å) show the presence of a hard active galactic nucleus (AGN) radiation field, while line ratio diagnostics are consistent with either AGN or star formation as the dominant ionizing source. GHZ9 is nitrogen-enriched (6–9.5 (N/O)⊙), carbon-poor (0.2–0.65 (C/O)⊙), metal-poor (Z = 0.01–0.1 Z⊙), and compact ( 106 pc), similarly to GN-z11, GHZ2, and recently discovered N-enhanced high redshift objects. We exploited the newly available JWST/NIRSpec and NIRCam data set to perform an independent analysis of the Chandra data confirming that GHZ9 is the most likely JWST source associated with X-ray emission at 0.5-7 keV. Assuming a spectral index = 2.3 (1.8), we estimate a black hole (BH) mass of 1.60 0.31 (0.48 0.09) 108M⊙, which is consistent either with Eddington-accretion onto heavy ( 106 M⊙) BH seeds formed at z=18, or super-Eddington accretion onto a light seed of 10 M⊙ at z = 25. The corresponding BH-to-stellar mass ratio MBH/Mstar= 0.330.22 (0.100.07), with a stringent limit 0.02, implies an accelerated growth of the BH mass with respect to the stellar mass. GHZ9 is the ideal target to constrain the early phases of AGN-galaxy coevolution with future multi-frequency observations.
@@ 1 Introduction
The James Webb Space Telescope (JWST) is revolutionizing our understanding of both galaxies and active galactic nuclei (AGN) in the high-redshift Universe. Several surveys have found a density of bright galaxies at , which is significantly larger than previously predicted (e.g., Castellano et al., 2022, 2023; Chemerynska et al., 2024; Finkelstein et al., 2024; McLeod et al., 2024). In addition, JWST NIRSpec (Jakobsen et al., 2022) observations have detected a higher-than-expected number of both broad-line AGN (BLAGN, e.g., Harikane et al., 2023; Maiolino et al., 2024a; Matthee et al., 2024) and narrow-line AGN (NLAGN, e.g., Chisholm et al., 2024; Curti et al., 2025) at high redshift, exceeding the predictions from extrapolated quasar (e.g., Niida et al., 2020; Shen et al., 2020) and AGN (e.g., Finkelstein & Bagley, 2022) luminosity functions, and from deep X-ray surveys (e.g., Giallongo et al., 2015, 2019).
@@
A comprehensive census of the AGN population at high redshifts is crucial for understanding the origin of the correlation between the physical properties of supermassive black holes (SMBHs) and their host galaxies, as found in the local Universe (e.g., Kormendy & Ho, 2013; Greene et al., 2020). In addition, a deeper understanding of the demographics and properties of distant AGN is necessary to assess their role in the reionization process (e.g., Giallongo et al., 2015; Madau et al., 2024; Dayal et al., 2025) and evaluate whether they contribute significantly to the ionizing UV emission from the bright galaxy population at z9, as suggested by recent studies (e.g., Castellano et al., 2024; Hegde et al., 2024; Maiolino et al., 2024a; Harikane et al., 2025). However, identifying the AGN population at high redshifts remains challenging.
@@
The demarcation lines between star-forming galaxies (SFGs) and AGN in the classic diagnostic diagrams (Baldwin et al., 1981; Veilleux & Osterbrock, 1987; Kauffmann et al., 2003) are less effective at high redshifts, with the two populations overlapping in the same regions (Kocevski et al., 2023; Maiolino et al., 2024a; Scholtz et al., 2025; Übler et al., 2023). New diagnostic diagrams have been calibrated (e.g., Calabrò et al., 2023; Hirschmann et al., 2023; Mazzolari et al., 2024b) and tested on a few z10 AGN candidates identified by JWST, such as GN-z11 (Maiolino et al., 2024b) and GHZ2 (Castellano et al., 2024).
@@
The discovery of AGN candidates beyond z=10, when the Universe was less than 450 Myr old, is pushing our exploration closer to the black hole (BH) seeding epoch, offering a unique opportunity to constrain the origin of SMBHs (Natarajan et al., 2024; Trinca et al., 2024b; Gordon et al., 2025; Hu et al., 2025; Jeon et al., 2025). As recently highlighted by Taylor et al. (2025a), this remote epoch represents a unique window to probe BH seeds, since by z 6 the BH mass function has largely lost memory of its initial seeding phase (see also, Valiante et al., 2018).
@@
The SMBH masses associated with the Chandra X-ray detections of the z10 objects UHZ1 (Goulding et al., 2023; Bogdán et al., 2024), and GHZ9 (Kovács et al., 2024) have prompted discussions on alternative BH seeding scenarios pathways beyond the standard light and heavy seed scenarios (Inayoshi et al., 2020), including models invoking accretion onto cosmological primordial BH seeds (Dayal, 2024; Matteri et al., 2025; Zhang et al., 2025; Ziparo et al., 2025).
@@
In this paper, we present a detailed analysis of the physical properties of GHZ9, based on the JWST/NIRSpec PRISM data presented in Napolitano et al. (2025), which confirmed the object at z=10.145. We examine the contribution of AGN and star formation to the spectrum of this source through rest-frame UV and optical diagnostic diagrams. We also take advantage of the newly available NIRSpec and NIRCam information on the sources in its vicinity to perform an independent analysis of the association with the Chandra X-ray emission, first presented by Kovács et al. (2024). GHZ9 provides a unique opportunity to test our understanding of AGN at high redshifts, as it is the most distant X-ray detected AGN known to date and benefits from several rest-frame optical and UV line detections thanks to JWST/NIRSpec.
@@ 2 NIRSpec observations and data analysis
GHZ9 (R.A. = 3.478756, decl. = -30.345520) was identified as a high-redshift candidate in the GLASS-JWST NIRCam field (Treu et al., 2022) by Castellano et al. (2023). It was observed using NIRSpec in the PRISM-CLEAR configuration as part of the Cycle 2 program GO-3073 (PI: Marco Castellano). The observation utilized three-shutter slits with a three-point nodding pattern for optimal background subtraction, with a total exposure time of 19,701 s over three separate visits.
@@
The detailed data reduction, as well as the GHZ9 spectrum and analysis (including redshift determination and line fitting), are presented in Napolitano et al. (2025). Briefly, data were processed using the standard calibration pipeline provided by STScI (version 1.13.4) and the Calibration Reference Data System mapping 1197, following the methodology of Arrabal Haro et al. (2023), which produces both 2D and 1D flux-calibrated spectra. To correct for potential slit-losses, the NIRSpec spectrum was calibrated against the most recent NIRCam broadband photometry (Merlin et al., 2024) by matching the continuum level. Additionally, since the source is magnified by the foreground Abell-2744 cluster, rest-frame quantities for GHZ9 were corrected for magnification ( = 1.36, Bergamini et al., 2023).
@@
The stellar masses of GHZ9 and of the two foreground interlopers ID = 29686 and ID = 29852 (see Fig. 1) were estimated using zphot (Fontana et al., 2000) as described by Santini et al. (2023), by fitting the observed HST and JWST photometry111The photometry of GHZ9 (ID = 29722) and of the two foreground sources is publicly available in the ASTRODEEP catalog (Merlin et al., 2024) https://vizier.cds.unistra.fr/viz-bin/VizieR?-source=J/A+A/691/A240 with Bruzual & Charlot (2003) templates, assuming delayed star formation histories (SFHs). The contribution from nebular continuum and line emission was included following Schaerer & de Barros (2009) and Castellano et al. (2014).
@@
We measured the half-light radius (re = 0.028′′) in the rest-frame UV using the same procedure adopted by Mascia et al. (2024), with the python software Galight222https://github.com/dartoon/galight (Ding et al., 2020). Assuming a Sérsic profile, with an axial ratio between and , and a Sérsic index of = 1, the fit was performed on the F150W NIRCam image. We visually inspected the result, finding no significant residuals in the luminosity profile.
@@
The spectroscopic redshift (zspec = 10.145 0.010) was determined from a weighted average of emission line centroids with signal-to-noise ratio (S/N) 5, calculated via direct integration. We measured the UV slope ( = -1.10 0.12) by fitting a power-law model () to the continuum flux at 1400–2600 Å rest-frame, after masking any potential emission features within the considered wavelength range. We employed emcee (Foreman-Mackey et al., 2013) for Markov chain Monte Carlo (MCMC) analysis.
@@
For emission lines with S/N 3, a Gaussian fit was applied to the continuum-subtracted flux using the specutils package from astropy (Astropy Collaboration et al., 2013) combined with emcee. Unresolved doublets and multiplets were modeled as single Gaussian profiles, while partially blended lines were fitted with double-Gaussian profiles (see Figure A.1 from Napolitano et al. (2025) for visualization).
@@ 3 Chandra X-ray observations
We analyzed all 101 publicly available Chandra X-ray observations of Abell-2744 (Chadayammuri et al., 2024), excluding ObsId=2212 because it was taken with a different CCD (ACIS-S instead of ACIS-I). We reprocessed all of the observations in the whole 0.5–7 keV band where Chandra is sensitive using standard CIAO tools and created a mosaic, with the deepest part having an exposure of 2.14 Ms. To ensure accurate astrometry, we matched sources detected in the 0.5–7 keV band to the Gaia DR2 catalog and found that a small translation of 0.27′′ was required, consistent with Kovács et al. (2024). GHZ9 has an average off-axis angle of 6.6′ from the deepest part of the X-ray field; the aperture radius including 90% of the encircled energy fraction (EEF) at 1.5 keV is 6′′. The net exposure time, accounting for vignetting, is 1.76 Ms. Given the presence of a nearby (4′′) bright X-ray star, we carefully accounted for its contamination.
@@
We extracted GHZ9 source counts from a circular region of 1′′ radius centered on the JWST position in the 0.5–7 keV band image, where a peak of emission is visible in X-rays (see Figure 1). This region includes only 13% of the total source counts due to the EEF. To compute the total number of counts, we considered both the nearby star contamination at the location of GHZ9 and the background contribution, evaluated from a nearby source-free circular region of 100 arcsec2. We also verified that the background is stable over a much larger region. In particular, we found that the background at the position of GHZ9 is consistent with the average background measured at the position of other X-ray sources detected outside the region affected by the diffuse emission from the cluster, whose peak (outskirts) is at 5.5 (2.8) arcmin from GHZ9. We measured 35.96.9 counts in the 0.5–7 keV band for GHZ9.
@@
We performed additional checks on the X-ray detection. As a first test, we ran the Chandra detection tool for point-like sources wavdetect in a 100100 arcsec2 region centered on GHZ9, after removing the star contamination. GHZ9 is detected with 23.27.6 counts in the observed-frame 0.5–7 keV band (3.3 significance). The flux and luminosity values discussed in the following would be scaled down by 35%. As a second check, we evaluated whether the shape of the X-ray continuum of GHZ9 differs from that of the nearby star by computing the ratio between the counts in the soft (S=0.5–2 keV) and hard (H=2–7 keV) bands for both objects. We found this test to be inconclusive due to the large uncertainties resulting in S/H=2.0 for GHZ9 and 1.2 for the star. As a final check, we divided the Chandra data set into two subsets with comparable exposure time. Unsurprisingly, the source is formally undetected in each individual subset. However, the counts measured within a 1″ radius circular region in both subsets are consistent with each other and with the expected signal given the reduced exposure time, thus excluding the possibility that the detection is due to a spurious fluctuation in the counts.
@@
The 0.5–7 keV counts for GHZ9 were then converted into flux and luminosity (both de-magnified) assuming a power-law model with two possible spectral indexes =1.8 (Piconcelli et al., 2005) and =2.3 (as in Kovács et al., 2024), which is comparable to the values measured in luminous quasars at z6–7.5 (Zappacosta et al., 2023). The assumption of the power-law, related to the main AGN emission process in X-rays (inverse Compton emission of accretion disk photons with energetic electrons in the corona), is required to transform the source count rate, which is a detector-dependent measure, into physical quantities, such as flux and luminosity. In the conversion process, we have considered the Chandra effective area corresponding to Cycle 24, during which most of the Abell 2744 observations were carried out. The observed-frame 2–10 keV flux is 1.310-16 erg cm-2s-1 if =2.3 (2.010-16 erg cm-2s-1 if =1.8). The corresponding rest-frame 2–10 keV luminosity is 3.81044 erg s-1 (1.81044 erg s-1). For the observed-frame 0.5–3 keV band, the flux measured, 2.210-16 erg cm-2s-1, is consistent with the value reported by Kovács et al. (2024), assuming the same photon index and magnification factor.